Find the equation of the right circular cone whose vertex is (1,−1,2), the axis is

(x−1 /2) =( y+1 /1) = (z−2 /−2)

and the semi-vertical angle is 45◦ .


1
Expert's answer
2021-06-09T10:20:33-0400

Let "P(x,y,z)" be any point on the cone. The direction ratio from the origin(axis) to point on the cone is "x-1, y+1, z-2" and the direction ratio of the axis are "2,1,-2."


Given the semi-vertical angle is 45◦


"\cos 45\degree=\dfrac{1}{\sqrt{2}}"

"=\dfrac{(x-1)(2)+( y+1)(1)+( z-2)(-2))}{\sqrt{(x-1)^2+(y+1)^2+(z-2)^2}\sqrt{(2)^2+(1)^2+(-2)^2}}"

"2(2x+y-2z+3)^2=9((x-1)^2+(y+1)^2+(z-2)^2)"




"8xy-16xz+8x^2+24x+2y^2+12y-8yz"

"+8z^2-24z+18=9x^2-18x+9"

"+9y^2+18y+9+9z^2-36z+36"



On simplifying this we get


"x^2+7y^2+z^2-8xy+8yz+16xz"

"-42x+6y-12z+36=0"